If Laplace transform method is used to solve the IVP: y(t) - 4 y(t) + 4y(t) = 4 cos2t, yO)= 2; y(O)=5 then the solution is
Answer #1

- کد ره )= 2 = (0) ره 2 - (515- ریوی : (تو ؟ ا رد الا - ای S1 :( کا (+2) sما 4 + 4 = Then . 4 both bansfor laplace take دورPartial fraction B 25-3 A 5-2 (5-272 (s-22 A = 2 B²1 0 2 S-2 ㅗ (3-222 +CS TO A § (5²4) 45 T (5-2)² S-2 (3²+4) (3-2)2 C=0, Dz-